Assuming 100% efficient energy conversion, how much water stored behind a 50
centimeter high hydroelectric dam would be required to charge the battery?
1
Expert's answer
2020-06-11T10:35:00-0400

Let us calculate the hydropower of the dam (see https://en.wikipedia.org/wiki/Hydropower#Mechanical_power)

"P = \eta \dot{m}g h" , where "\eta" is the efficiency, "\dot{m}" is the mass flow rate, "g" is the acceleration due to gravity and "h" is the height.

If we calculate the energy provided by dam, we should multiply the formula by time, so we'll get

"E=\eta mgh."

There are many different types of batteries. AAA batteries have typical capacity of ~ 1 Wh = "3.6\cdot10^3\,"J.

Therefore, "m = \dfrac{E}{\eta gh} = \dfrac{3.6\cdot10^3\,\mathrm{J}}{1.00\cdot9.81\,\mathrm{N/kg} \cdot0.50\,\mathrm{m}} = 734\,\mathrm{kg}."

If we consider D batteries, we'll read they have capacities of ~ 10 Wh (see https://en.wikipedia.org/wiki/D_battery). Therefore,

"m = \dfrac{E}{\eta gh} = \dfrac{3.6\cdot10^4\,\mathrm{J}}{1.00\cdot9.81\,\mathrm{N/kg} \cdot0.50\,\mathrm{m}} \approx 7340\,\mathrm{kg}."